OpenStudy (anonymous):

is 2016 the least common multiple of 32,56,72 and show how u got the answer please

OpenStudy (anonymous):

There are many ways to find the LCM. Here's one way using prime factorizations. 32=2x2x2x2x2=2^5 56=2x2x2x7=2^3x7 72=2x2x2x3x3=2^3x3^2 LCM = 2^5x3^2x7 = 2016